### v1

# Detection and Reconstruction of a Random Hypergraph from Noisy Graph Projection

Abstract
For a $d$-uniform random hypergraph on $n$ vertices in which hyperedges are included i.i.d. so that the average degree in the hypergraph is $n^{\delta+o(1)}$, the projection of such a hypergraph is a graph on the same $n$ vertices where an edge connects two vertices if and only if they belong to a same hyperedge. In this work, we study the inference problem where the observation is a noisy version of the graph projection where each edge in the projection is kept with probability $p=n^{-1+\alpha+o(1)}$ and each edge not in the projection is added with probability $q=n^{-1+\beta+o(1)}$. For all constant $d$, we establish sharp thresholds for both detection (distinguishing the noisy projection from an Erdős-Rényi random graph with edge density $q$) and reconstruction (estimating the original hypergraph). Notably, our results reveal a detection-reconstruction gap phenomenon in this problem. Our work also answers a problem raised in [BGPY25].
1 Introduction
In this paper, we address the problem of reconstructing the multi-vertex associations of a hypergraph from its projected graph, a graph in which two vertices are connected if and only if they belong to the same hyperedge. More precisely, we pose the reconstruction task as follows: for any constant $d\in\mathbb{N}$, we denote by $\operatorname{U}=\operatorname{U}_{n,d}$ the collection containing all subsets of $[n]=\{1,\ldots,n\}$ with cardinality $d$.
Definition 1.1.
Define $\mathcal{H}$ to be the random $d$-uniform hypergraph on $[n]$ generated by including each $\Psi\in\operatorname{U}_{n,d}$ in $\mathcal{H}$ with probability $s=n^{-(d-1)+\delta+o(1)}$ for some constant $10\leq\delta<$ $1$For any two positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write equivalently $a_{n}=O(b_{n})$, $b_{n}=\Omega(a_{n})$, $a_{n}\lesssim b_{n}$ and $b_{n}\gtrsim a_{n}$ if there exists a positive absolute constant $c$ such that $a_{n}/b_{n}\leq c$ holds for all $n$. We write $a_{n}=o(b_{n})$, $b_{n}=\omega(a_{n})$, $a_{n}\ll b_{n}$, and $b_{n}\gg a_{n}$ if $a_{n}/b_{n}\to 0$ as $n\to\infty$. We write $a_{n}=\Theta(b_{n})$ if both $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$ hold. Without further specification, all asymptotics are taken with respect to $n\to\infty$.. We denote the law of $\mathcal{H}$ by $\mu=\mu_{n,d,\delta}$. For simplicity, we will also use $\mathcal{H}$ to denote the set of hyperedges and $|\mathcal{H}|$ the cardinality of $\mathcal{H}$. In addition, for any $d$-uniform hypergraph $\mathcal{H}$, define $\mathcal{P}(\mathcal{H})$ to be the projection graph on $[n]$ where for all unordered pairs $(i,j)$ with $i\neq j$
$(i,j)\in\mathcal{P}(\mathcal{H})\Longleftrightarrow\exists\Psi\in\mathcal{H}\mbox{ such that }i,j\in\Psi\,.$ (1.1)
Consider the random graph $A\triangleq A(\mathcal{H})$ on $[n]$ obtained according to the following rule (denoted by $\mathbb{P}=\mathbb{P}_{n,d,\delta,\alpha,\beta}$): we first sample $\mathcal{H}\sim\mu$. Then (conditioned on $\mathcal{H}$)
  • for each $(i,j)\in\mathcal{P}(\mathcal{H})$, we independently include $(i,j)$ in $A$ with probability $p=n^{-1+\alpha+o(1)}\leq 1$ for some $0\leq\alpha\leq 1$;
  • for each $(i,j)\not\in\mathcal{P}(\mathcal{H})$, we independently include $(i,j)$ in $A$ with probability $q=n^{-1+\beta+o(1)}\leq p$.
We require the condition $q\leq p$, reflecting the natural expectation that elements of $\mathcal{P}(\mathcal{H})$ tend to appear in the graph $A$. Inspired by real-world social networks such as scientific co-authorship networks [New04], graphs with group structure (e.g., random intersection graph [BBN20, vdHKV21] and hypergraph stochastic block model [GD17, ACKZ15, SZ22, GJ23, DWar, GP23, GP24]) have garnered significant attention in the research community. Among them, the model of noisy random hypergraph projection serves as a natural toy model to study how to infer the original group structure given only pairwise vertex information. For a more comprehensive overview on this model and its various applications in real networks, we refer the readers to [BGPY25]. Two natural inference problems regarding this model are as follows: (1) the detection problem, i.e., testing $\mathbb{P}$ against $\mathbb{Q}=\mathbb{Q}_{n,q}$, where $\mathbb{Q}$ is the law of an Erdős-Rényi graph on $[n]$ with edge density $q=n^{-1+\beta+o(1)}$; (2) the reconstruction problem, i.e., estimating the original hypergraph $\mathcal{H}$ according to the observation $A(\mathcal{H})$.
The problem of reconstruction from the projected graph has been extensively studied in recent years, both theoretically and empirically. For instance, [WK24] assumes access to a sample of a hypergraph from the underlying distribution and proposes a scoring method to select hyperedges based on their similarity to the sampled hypergraph. Another Bayesian algorithmic approach (i.e., to sample from the posterior distribution given the projected graph) was proposed in [YPP21, LYA23]. In addition, in [CFLL24] a large foundational model is used to recover a weighted hypergraph from a sample of its hyperedges, where each hyperedge is assigned a probability proportional to its weight. In a series of recent progress [BGP24, BGPY25], the sharp threshold for reconstructing the hypergraph $\mathcal{H}$ based on $A(\mathcal{H})$ in the noiseless setting $p=1,q=0$ was determined. In particular, it was shown in [BGPY25] that when $\delta>\tfrac{d-1}{d+1}$ it is information theoretically impossible to reconstruct a positive fraction of $\mathcal{H}$ and when $\delta<\tfrac{d-1}{d+1}$ there exists an efficient algorithm that reconstructs a $1-o(1)$ fraction of $\mathcal{H}$. They further raised the question of whether an analogous reconstruction threshold can be established in the noisy setting. Our main result can be informally summarized as follows:
Theorem 1.2.
  1. (1)
    (Detection) Suppose $\beta>2(\delta+\alpha)-1$. Then it is information theoretically impossible to distinguish $\mathbb{P}$ from $\mathbb{Q}$. Conversely, if $\beta<2(\delta+\alpha)-1$. Then there exists an efficient algorithm to distinguish $\mathbb{P}$ from $\mathbb{Q}$.
  2. (2)
    (Partial reconstruction) Suppose either $p=o(1)$ or $\delta>\tfrac{d-1}{d+1}$ or $\beta>\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$. Then it is informationally impossible to reconstruct a positive fraction of $\mathcal{H}$. Conversely, if $p=\Omega(1),\delta<\tfrac{d-1}{d+1}$ and $\beta<\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$, then there exists an efficient algorithm that successfully reconstructs a positive fraction of $\mathcal{H}$.
  3. (3)
    (Almost-exact reconstruction) Suppose either $p=1-\Omega(1)$, $\delta>\tfrac{d-1}{d+1}$ or $\beta>\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$. Then it is informationally impossible to reconstruct a $1-o(1)$ fraction of $\mathcal{H}$. Conversely, if $p=1-o(1),\delta<\tfrac{d-1}{d+1}$ and $\beta<\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$, then there exists an efficient algorithm that successfully reconstructs a $1-o(1)$ fraction of $\mathcal{H}$.
Notably, our result reveals a detection-reconstruction gap phenomenon in this problem where when $p=\Theta(1)$ (which implies that $\alpha=1$), detection is possible for all $\beta,\delta\in(0,1)$ but (partial) reconstruction is possible if and only if $\delta<\tfrac{d-1}{d+1}$ and $\beta<\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$. See Figure 1 for a summary of the regions where detection/reconstruction is possible. The reader may wonder whether the detection-recovery gap could be eliminated by choosing an alternative null distribution whose edge density matched that of the planted model $\mathbb{P}$. We show in Appendix A that this adjustment closes the gap partially but not entirely. In particular, detection remains possible at least whenever $\delta,\beta<\tfrac{d-1}{d+1}$, noting that for $\delta<\tfrac{d-1}{d+1}$ we always have $\tfrac{d-1}{d+1}>\tfrac{d-2}{d}+\tfrac{2\delta}{d(d-1)}$.
Figure 1: Summary of the detection-reconstruction gap when $p=\Theta(1)$. Red regime $I$: reconstruction of $\mathcal{H}$ is possible; Blue regime $II$: detection with a better null distribution $\widetilde{\mathbb{Q}}$ (that matches the average edge-density) is possible, but reconstruction of $\mathcal{H}$ is impossible; Yellow regime $III$: detection with the simple null distribution $\mathbb{Q}$ is possible, but detection with a better null distribution $\widetilde{\mathbb{Q}}$ is unknown.
1.1 The detection threshold
In this subsection, we state precisely the results that establish Item (1) of Theorem 1.2.
Definition 1.3.
For two sequences of probability measures $\mathbb{P}=\mathbb{P}_{n}$ and $\mathbb{Q}=\mathbb{Q}_{n}$ on the same sample $X=X_{n}$. A test statistic $f=f(X)\in\{0,1\}$ is said to achieve
  • Strong detection, if $\mathbb{P}(f(X)=0)+\mathbb{Q}(f(X)=1)=o(1)$ as $n\to\infty$.
  • Weak detection, if $\mathbb{P}(f(X)=0)+\mathbb{Q}(f(X)=1)=1-\Omega(1)$ as $n\to\infty$.
We first show that when $\beta<2(\delta+\alpha)-1$ there exists a test statistic that is efficient and achieves strong detection. Consider the following edge-counting statistic
${}f=f(A)=\sum_{1\leq i<j\leq n}\mathbf{1}_{\{(i,j)\in A\}}\,.$ (1.2)
It is clear that $f$ can be computed in time $O(n^{2})$.
Proposition 1.4.
Suppose that $\beta<2(\delta+\alpha)-1$. Then we have
$\mathbb{P}\Big{(}f(A)>q\tbinom{n}{2}+\tfrac{1}{2}ps\cdot\tbinom{d}{2}\tbinom{n}{d}\Big{)}=1-o(1)\,;$ (1.3)
$\mathbb{Q}\Big{(}f(A)<q\tbinom{n}{2}+\tfrac{1}{2}ps\cdot\tbinom{d}{2}\tbinom{n}{d}\Big{)}=1-o(1)\,.$
In particular, the thresholding statistic on $f(A)$ achieves strong detection between $\mathbb{P}$ and $\mathbb{Q}$.
The proof of Proposition 1.4 is incorporated in Section 2. We now turn to showing that whenever $\beta>2(\delta+\alpha)-1$, no test, strong or weak, can reliably distinguish $\mathbb{P}$ from $\mathbb{Q}$. It is well-known that the testing error is governed by the total variation distance between the null and alternative distributions. Our next result establishes the desired bound on this total variation distance in the regime $\beta>2(\delta+\alpha)-1$.
Proposition 1.5.
When $\beta>2(\delta+\alpha)-1$, we have $\operatorname{TV}(\mathbb{P},\mathbb{Q})=o(1)$. Thus, it is informationally impossible to achieve weak detection between $\mathbb{P}$ and $\mathbb{Q}$.
The proof of Proposition 1.5 is incorporated in Section 3. Our strategy is as follows: first we combine the standard Pinsker’s inequality and replica’s trick to bound $\operatorname{TV}(\mathbb{P},\mathbb{Q})$ by $E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))$ where $\mathcal{H},\mathcal{H}^{\prime}$ are independent copies from $\mu$. Then by an application of Talagrand’s concentration inequality we can argue that $|E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))|$ tightly concentrates around its mean, which yields the desired total variation bound.
1.2 The reconstruction threshold
In this subsection, we state precisely the results that confirm Items (2) and (3) of Theorem 1.2.
Definition 1.6.
Recall the definition of $s$ in 1.1. A test statistic ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ is said to achieve
  • Partial reconstruction, if $\mathbb{E}_{\mathbb{P}}\big{[}|\mathcal{H}\triangle\widehat{\mathcal{H}}|\big{]}=(1-\Omega(1))\cdot s\tbinom{n}{d}$ as $n\to\infty$.
  • Almost-exact reconstruction, if $\mathbb{E}_{\mathbb{P}}\big{[}|\mathcal{H}\triangle\widehat{\mathcal{H}}|\big{]}=o(1)\cdot s\tbinom{n}{d}$ as $n\to\infty$.
We first show the positive results on partial and almost-exact reconstruction. Consider the “clique” estimator
^:={${}\widehat{\mathcal{H}}:=\big{\{}\Psi\in\operatorname{U}:A_{i,j}=1\mbox{ for all }i,j\in\Psi\big{\}}\,.$ (1.4)
Proposition 1.7.
Suppose $\delta<\tfrac{d-1}{d+1},\beta<\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$ and $p=1-o(1)$. We have
$\mathbb{E}_{\mathbb{P}}\Big{[}\big{|}\widehat{\mathcal{H}}\triangle\mathcal{H}\big{|}\Big{]}=o(1)\cdot s\tbinom{n}{d}\,.$
In addition, when $\delta<\tfrac{d-1}{d+1},\beta<\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$ and $p=\Omega(1)$, we have
$\mathbb{E}_{\mathbb{P}}\Big{[}\big{|}\widehat{\mathcal{H}}\triangle\mathcal{H}\big{|}\Big{]}=(1-\Omega(1))\cdot s\tbinom{n}{d}\,.$
The proof of Proposition 1.7 is incorporated in Section 4. Next we show the impossibility of weak reconstruction. Note that when $\delta>\tfrac{d-1}{d+1}$, using [BGPY25, Theorem 1.1] it is informationally impossible to achieve weak reconstruction even when $p=1,q=0$. Thus, without loss of generality, we assume that $\delta\leq\tfrac{d-1}{d+1}$.
Proposition 1.8.
  • (1)
    Suppose that $p=o(1)$. Then for any estimator ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ we have
    $\mathbb{E}_{\mathcal{H},A\sim\mathbb{P}}\Big{[}\big{|}\widehat{\mathcal{H}}(A)\triangle\mathcal{H}\big{|}\Big{]}=(1-o(1))\cdot s\tbinom{n}{d}$
    even when $q=0$.
  • (2)
    Suppose $q=n^{-1+\beta+o(1)}$ where $\beta>\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$. Then for any estimator ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ we have
    $\mathbb{E}_{\mathcal{H},A\sim\mathbb{P}}\Big{[}\big{|}\widehat{\mathcal{H}}(A)\triangle\mathcal{H}\big{|}\Big{]}=(1-o(1))\cdot s\tbinom{n}{d}$
    even when $p=1$.
The proof of Proposition 1.8 is given in Section 5. Item (1) follows from a suitable generalization of Fano’s inequality, while Item (2) is established via a Bayesian argument with appropriate truncations.
We then show the impossibility of strong reconstruction. In light of Proposition 1.8, it remains only to prove the following result, whose proof appears in Section 6.
Proposition 1.9.
Suppose that $p=1-\Omega(1)$. Then for any estimator ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ we have
$\mathbb{E}_{\mathcal{H},A\sim\mathbb{P}}\big{[}|\widehat{\mathcal{H}}(A)\triangle\mathcal{H}|\big{]}=\Omega(1)\cdot s\tbinom{n}{d}$
even when $q=0$.
2 Proof of Proposition 1.4
Recall (1.2). Under $\mathbb{Q}$, $f(A)$ is the sum of $\tbinom{n}{2}$ i.i.d. random variables with law $\operatorname{Ber}(q)$. It is clear that $\mathbb{E}_{\mathbb{Q}}[f(A)]=q\tbinom{n}{2}$ and $\operatorname{Var}_{\mathbb{Q}}[f(A)]\leq q\tbinom{n}{2}$. By Chebyshev’s inequality, we get (also recall that $q=n^{-1+\beta+o(1)}$, $p=n^{-1+\alpha+o(1)}$ and $s=n^{-(d-1)+\delta+o(1)}$)
$\mathbb{Q}\Big{(}f(A)\geq q\tbinom{n}{2}+\tfrac{1}{2}ps\cdot\tbinom{d}{2}\tbinom{n}{d}\Big{)}\leq\frac{\operatorname{Var}_{\mathbb{Q}}[f(A)]}{\big{[}\tfrac{1}{2}ps\cdot\tbinom{d}{2}\tbinom{n}{d}\big{]}^{2}}=O\big{(}n^{1+\beta-2(\alpha+\delta+o(1))}\big{)}=o(1)\,,$
where the last equality follows from $\beta<2(\alpha+\delta)-1$. Now we focus on the behavior of $f(A)$ under $\mathbb{P}$. Note that conditioning on $\mathcal{H}$, we have (the binomial variables are independent)
[…]
Thus, we first show
$\mathbb{P}\Big{(}0.9s\tbinom{d}{2}\tbinom{n}{d}\leq|E(\mathcal{P}(\mathcal{H}))|\leq 1.1s\tbinom{d}{2}\tbinom{n}{d}\Big{)}=1-o(1)\,.$ (2.2)
Let $M_{k}(\mathcal{H})$ be the set of $(i,j)\in\mathcal{P}(\mathcal{H})$ such that there exists at least $k$ hyperedges in $\mathcal{H}$ that contain $(i,j)$. In addition, define
${}\mathcal{E}=\mathcal{E}_{1}\cap\mathcal{E}_{2}\,,$ (2.3)
where
${}\mathcal{E}_{1}=\Big{\{}|\mathcal{H}|=(1+o(1))s\tbinom{n}{d}\Big{\}}$ (2.4)
and
${}\mathcal{E}_{2}=\Big{\{}\sum_{k\geq 2}k|M_{k}(\mathcal{H})|=o(1)\cdot s\tbinom{n}{d}\Big{\}}\,.$ (2.5)
Clearly we have $\mathbb{P}_{\mathcal{H}\sim\mu}(\mathcal{E}_{1})=1-o(1)$. In addition, note that $(i,j)\in M_{k}(\mathcal{H})$ implies that there exist $k$ different $\Psi_{1},\ldots,\Psi_{k}\in\operatorname{U}$ with $\Psi_{l}\in\mathcal{H}$ and $(i,j)\in E(\mathcal{P}(\Psi_{l}))$ for $1\leq l\leq k$. Thus, by a simple union bound we get that (note that the enumeration of $k$ different $\Psi_{1},\ldots,\Psi_{k}\in\operatorname{U}$ with $(i,j)\in E(\mathcal{P}(\Psi_{l}))$ is bounded by $n^{(d-2)k}$)
$\mathbb{P}_{\mathcal{H}\sim\mu}\big{(}(i,j)\in M_{k}(\mathcal{H})\big{)}\leq n^{(d-2)k}\cdot s^{k}\leq n^{(-1+\delta+o(1))k}\,.$
This implies that
$\mathbb{E}_{\mathcal{H}\sim\mu}\Big{[}\sum_{k\geq 2}k|M_{k}(\mathcal{H})|\Big{]}\leq\sum_{k\geq 2}kn^{2}\cdot n^{(-1+\delta+o(1))k}=(1+o(1))n^{2\delta+o(1)}=o(1)\cdot s\tbinom{n}{d}\,.$
Thus, from a simple Markov inequality we get that $\mathbb{P}_{\mathcal{H}\sim\mu}(\mathcal{E}_{2})=1-o(1)$. Thus, we have
${}\mathbb{P}_{\mathcal{H}\sim\mu}(\mathcal{E})=1-o(1)\,.$ (2.6)
Also, under $\mathcal{E}$ we get from the deterministic bound that
$|\mathcal{P}(\mathcal{H})|\leq\tbinom{d}{2}|\mathcal{H}|\leq 1.1s\tbinom{d}{2}\tbinom{n}{d}$
and
$|E(\mathcal{P}(\mathcal{H}))|\geq\tbinom{d}{2}|\mathcal{H}|-\sum_{k\geq 2}k|M_{k}(\mathcal{H})|\geq 0.9s\tbinom{d}{2}\tbinom{n}{d}\,,$
which yields (2.2). Now conditioning on $\mathcal{P}(\mathcal{H})$ satisfying (2.2), by (2.1) and a standard binomial tail estimate, we obtain that with probability $1-o(1)$ under $\mathbb{P}$
$|E(A(\mathcal{H}))|\geq q\tbinom{n}{2}+0.9(p-q)s\tbinom{d}{2}\tbinom{n}{d}-\sqrt{(ps\tbinom{d}{2}\tbinom{n}{d}+q\tbinom{n}{2})\log n}\geq q\tbinom{n}{2}+\tfrac{1}{2}ps\tbinom{d}{2}\tbinom{n}{d}\,,$
where in the last inequality we used $n^{\frac{1+\beta}{2}}<n^{\delta+\alpha-\Omega(1)}$. As a result, (1.3) holds.
3 Proof of Proposition 1.5
Let $L(A)=\tfrac{\mathbb{P}(A)}{\mathbb{Q}(A)}$ denote the likelihood ratio. By Pinsker’s inequality we have
$\operatorname{TV}(\mathbb{P},\mathbb{Q})\leq\mathbb{E}_{\mathbb{Q}}\big{[}L(A)^{2}\big{]}-1\,.$
Thus, to prove Proposition 1.5, it suffices to show the following result.
Lemma 3.1.
When $\beta>2(\delta+\alpha)-1$, we have
${}\mathbb{E}_{\mathbb{Q}}\big{[}L(A)^{2}\big{]}=1+o(1)\,.$ (3.1)
To prove Lemma 3.1, we first bound $\mathbb{E}_{\mathbb{Q}}\big{[}L(A)^{2}\big{]}$ by the following lemma.
Lemma 3.2.
We have
${}\mathbb{E}_{\mathbb{Q}}\Big{[}L(A)^{2}\Big{]}\leq\mathbb{E}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu\otimes\mu}\Bigg{[}\exp\Big{(}n^{-1-\beta+2\alpha+o(1)}\cdot\big{|}E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))\big{|}\Big{)}\Bigg{]}\,.$ (3.2)
Proof.
Note that
$L(A)=\tfrac{\mathbb{P}(A)}{\mathbb{Q}(A)}=\mathbb{E}_{\mathcal{H}\sim\mu}\Big{[}\tfrac{\mathbb{P}(A\mid\mathcal{H})}{\mathbb{Q}(A)}\Big{]}=\mathbb{E}_{\mathcal{H}\sim\mu}\Big{[}\prod_{1\leq i<j\leq n}\tfrac{\mathbb{P}(A_{i,j}\mid\mathcal{H})}{\mathbb{Q}(A_{i,j})}\Big{]}=\mathbb{E}_{\mathcal{H}\sim\mu}\Big{[}\prod_{(i,j)\in E(\mathcal{P}(\mathcal{H}))}\ell(A_{i,j})\Big{]}\,,$
where
$\ell(A_{i,j})=\mathbf{1}_{\{A_{i,j}=1\}}\cdot\tfrac{n^{-1+\alpha+o(1)}}{n^{-1+\beta+o(1)}}+\mathbf{1}_{\{A_{i,j}=0\}}\cdot\tfrac{1-n^{-1+\alpha+o(1)}}{1-n^{-1+\beta+o(1)}}\,.$
We then have (note that $\mathbb{E}_{\mathbb{Q}}[\ell(A_{i,j})]=1$)
$\mathbb{E}_{\mathbb{Q}}\Big{[}L(A)^{2}\Big{]}$ $=\mathbb{E}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu\otimes\mu}\Bigg{\{}\mathbb{E}_{A\sim\mathbb{Q}}\Big{[}\tfrac{\mathbb{P}(A\mid\mathcal{H})}{\mathbb{Q}(A)}\cdot\tfrac{\mathbb{P}(A\mid\mathcal{H}^{\prime})}{\mathbb{Q}(A)}\Big{]}\Bigg{\}}$
$=\mathbb{E}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu\otimes\mu}\Bigg{\{}\mathbb{E}_{A\sim\mathbb{Q}}\Big{[}\prod_{(i,j)\in E(\mathcal{P}(\mathcal{H}))}\ell(A_{i,j})\cdot\prod_{(i,j)\in E(\mathcal{P}(\mathcal{H}^{\prime}))}\ell(A_{i,j})\Big{]}\Bigg{\}}$
$=\mathbb{E}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu\otimes\mu}\Bigg{\{}\prod_{(i,j)\in E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))}\mathbb{E}_{A\sim\mathbb{Q}}\Big{[}\ell(A_{i,j})^{2}\Big{]}\Bigg{\}}\,.$ (3.3)
Finally, from straightforward calculation we have
$\mathbb{E}_{A\sim\mathbb{Q}}\Big{[}\ell(A_{i,j})^{2}\Big{]}$ $=\Big{(}\tfrac{n^{-1+\alpha+o(1)}}{n^{-1+\beta+o(1)}}\Big{)}^{2}\cdot n^{-1+\beta+o(1)}+\Big{(}\tfrac{1-n^{-1+\alpha+o(1)}}{1-n^{-1+\beta+o(1)}}\Big{)}^{2}\cdot(1-n^{-1+\beta+o(1)})$
$\leq n^{-1-\beta+2\alpha+o(1)}+1\leq\,\exp\big{(}n^{-1-\beta+2\alpha+o(1)}\big{)}\,,$
where the second line follows from the assumption $\beta\leq\alpha\leq 1$. Plugging the above estimation into (3.3), we get (3.2). ∎
Based on Lemma 3.2, it suffices to bound the tail probability of $|E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))|$, as incorporated in the following lemma.
Lemma 3.3.
We have
${}\mathbb{P}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu\otimes\mu}\Big{(}\big{|}E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))\big{|}\geq n^{2\delta+o(1)}+tn^{\delta+o(1)}\Big{)}\leq\exp\big{(}-\Omega(1)\cdot t^{2}\big{)}\,.$ (3.4)
To prove Lemma 3.3, we begin by recalling the relevant definitions and notation for Talagrand’s concentration inequality.
Definition 3.4.
Let $\Omega=\prod_{i=1}^{N}\Omega_{i}$ be a product space. For a function $h:\Omega\to\mathbb{R}$, we call $h$ to be Lipschitz, if $|h(x)-h(y)|\leq 1$ whenever $x,y$ differ in at most one coordinate. Furthermore, for an increasing function $f:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$, we say that $h$ is $f$-certifiable if, whenever $h(x)\geq s$, there exists $I\subset\{1,\ldots,N\}$ with $|I|\leq f(s)$ such that all $y\in\Omega$ that agree with $x$ on the coordinates in $I$ satisfy $h(y)\geq s$.
A special case of Talagrand’s concentration inequality can be stated as follows.
Theorem 3.5.
Let $\mathbb{P}$ be any product measure on a sample space $\Omega$, and let $h$ be a Lipschitz function that is $f$-certifiable. Writing $X=h(\omega)\,(\omega\in\Omega)$, for any $b,t>0$ we have:
$\mathbb{P}\big{(}X\leq b-t\sqrt{f(b)}\big{)}\cdot\mathbb{P}\big{(}X\geq b\big{)}\leq e^{-t^{2}/4}\,.$
The proof of inequality can be found in, e.g., [AS16, Section 7.7]. We now proceed to the proof of Lemma 3.3.
Proof of Lemma 3.3.
For our purpose, we consider $\Omega=\{0,1\}^{2\binom{n}{d}}$ and $\mathbb{P}$ be the product measure of $2\binom{n}{d}$ Bernoulli variables with parameter $s=n^{-(d-1)+\delta+o(1)}$. Let
$X=h(\mathcal{H},\mathcal{H}^{\prime})=\big{|}E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))\big{|}/\tbinom{d}{2},$
where we view $(\mathcal{H},\mathcal{H}^{\prime})$ as a vector in $\{0,1\}^{2\binom{n}{d}}$. It is straightforward to verify that $h$ is Lipschitz. Furthermore, we claim that the function $h$ defined above is $f$-certifiable, where $f(s)=2\tbinom{d}{2}s\leq d^{2}s$. In fact, if $h(\mathcal{H},\mathcal{H}^{\prime})\geq s$ for some $(\mathcal{H},\mathcal{H}^{\prime})\in\Omega$, then there exists $E=\{(i,j):1\leq i<j\leq n\}$ such that for all $e\in E$ there exist $\zeta_{e},\zeta_{e}^{\prime}\in\tbinom{[n]}{d}$ with $\mathcal{H}(\zeta_{e})=\mathcal{H}^{\prime}(\zeta^{\prime}_{e})=1$. Denote $I=\{\zeta_{e}:e\in E\}\cap\{\zeta^{\prime}_{e}:e\in E\}$. Then for all (~,~)$(\widetilde{\mathcal{H}},\widetilde{\mathcal{H}}^{\prime})\in\{0,1\}^{2\binom{n}{d}}$ that agrees with $(\mathcal{H},\mathcal{H}^{\prime})$ on $I$, we have $e\in E(\mathcal{P}(\widetilde{\mathcal{H}}))\cap E(\mathcal{P}(\widetilde{\mathcal{H}}^{\prime}))$ and thus $h(\widetilde{\mathcal{H}},\widetilde{\mathcal{H}}^{\prime})\geq s$. As a result, Theorem 3.5 yields that for any $b,t\geq 0$,
$\mathbb{P}\big{(}X\leq b-td\sqrt{2b}\big{)}\cdot\mathbb{P}(X\geq b)\leq e^{-t^{2}/4}\,.$
Taking $b=\operatorname{Median}(X)$ to be the median of $X$, we then get that
$\mathbb{P}\Big{(}X\leq\operatorname{Median}(X)-t\Big{)}\leq 2e^{-t^{2}/4d\operatorname{Median}(X)}\,.$
In addition, taking $b=\operatorname{Median}(X)+r$ and $t=r/\sqrt{b}$ we get that
$\mathbb{P}\Big{(}X\geq\operatorname{Median}(X)+r\Big{)}\leq 2e^{-r^{2}/4d(\operatorname{Median}(X)+r)}\leq 2e^{-r/8d}+2e^{-r^{2}/8d\operatorname{Median}(X)}\,.$
Thus we have
$\Big{|}\mathbb{E}[X]-\operatorname{Median}(X)\Big{|}$ $\leq\mathbb{E}\Big{[}\big{|}X-\operatorname{Median}(X)\big{|}\Big{]}=\int_{0}^{\infty}\mathbb{P}\Big{(}\big{|}X-\operatorname{Median}(X)\big{|}\geq t\Big{)}\mathrm{d}t$
$\leq\int_{0}^{\infty}4\exp\Big{(}-\tfrac{t^{2}}{8d\operatorname{Median}(X)}\Big{)}+4\exp(-t)\mathrm{d}t\leq 10d\sqrt{\operatorname{Median}(X)}\,.$ (3.5)
By the independence of $\mathcal{H}$ and $\mathcal{H}^{\prime}$, we can easily get
$\mathbb{E}[X]=\tbinom{n}{2}\big{(}\tbinom{n-2}{d-2}s\big{)}^{2}/\tbinom{d}{2}=n^{2\delta+o(1)}\,.$
Thus, we have (also note that $\mathbb{E}[X]=n^{2\delta+o(1)}$ and (3.5) imply that $\operatorname{Median}(X)=\mathbb{E}[X]+\Omega(1)\cdot\sqrt{\mathbb{E}[X]}$)
${}\mathbb{P}\Big{(}X\geq\mathbb{E}[X]+t\sqrt{\mathbb{E}[X]}\Big{)}\leq 2e^{-\Omega(1)\cdot t^{2}}\,.$ (3.6)
Plugging $\mathbb{E}[X]=n^{2\delta+o(1)}$ into (3.6) yields the desired result. ∎
Now we can complete the proof of Lemma 3.1.
Proof of Lemma 3.1.
Let $Y(\mathcal{H},\mathcal{H}^{\prime})=\big{|}E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))\big{|}$ and recall that $1+\beta>2(\delta+\alpha)$. Applying Lemma 3.3, we obtain the following estimate
$\quad\ \mathbb{E}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu\otimes\mu}\Big{[}\exp\big{(}n^{-1-\beta+2\alpha+o(1)}\cdot Y\big{)}\cdot\mathbbm{1}\big{\{Y\geq 2n^{2\delta+o(1)}\big{\}}}\Big{]}$
$=\int_{2n^{2\delta+o(1)}}^{\infty}n^{-1-\beta+2\alpha+o(1)}\exp\big{(}n^{-1-\beta+2\alpha+o(1)}\cdot y\big{)}\cdot\mathbb{P}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu\otimes\mu}(Y\geq y)\mathrm{d}y$
${\leq}\int_{n^{\delta+o(1)}}^{\infty}\exp\Big{(}n^{-1-\beta+2\alpha+o(1)}\big{(}n^{2\delta+o(1)}+tn^{\delta+o(1)}\big{)}\Big{)}\exp(-ct^{2})\mathrm{d}t$
$=o\Big{(}\exp\big{(}n^{-1-\beta+2(\alpha+\delta)+o(1)}\big{)}\Big{)}=o(1).$
Then, we have (recall that $1+\beta=2(\alpha+\delta)+\Omega(1)$)
$\mathbb{E}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu\otimes\mu}\Big{[}\exp\big{(}n^{-1-\beta+2\alpha+o(1)}\cdot Y\big{)}\Big{]}\leq\exp\big{(}2\cdot n^{-1-\beta+2(\alpha+\delta)+o(1)}\big{)}+o(1)=1+o(1).$
Combining this bound with Lemma 3.2 gives (3.1). ∎
4 Proof of Proposition 1.7
Firstly, note that for all $\Psi\subset[n]$ with $|\Psi|=d$ we have
$\mathbb{P}(\Psi\in\widehat{\mathcal{H}}\cap\mathcal{H})=p^{\binom{d}{2}}\cdot s=\Theta(1)\cdot s=\Theta(1)\cdot\mathbb{P}(\Psi\in\mathcal{H})\mbox{ when }p=\Omega(1)\,,$
$\mathbb{P}(\Psi\in\widehat{\mathcal{H}}\cap\mathcal{H})=p^{\binom{d}{2}}\cdot s=(1-o(1))\cdot s=(1-o(1))\cdot\mathbb{P}(\Psi\in\mathcal{H})\mbox{ when }p=1-o(1)\,.$
This implies that
$\mathbb{E}\left[|\mathcal{H}\setminus\widehat{\mathcal{H}}|\right]=(1-\Omega(1))\cdot s\tbinom{n}{d}\text{ when }p=\Omega(1)\,,$
$\mathbb{E}\left[|\mathcal{H}\setminus\widehat{\mathcal{H}}|\right]=o(1)\cdot s\tbinom{n}{d}\text{ when }p=1-o(1)\,.$
Thus, it suffices to show that when $\delta<\tfrac{d-1}{d+1}$ and $\beta<\tfrac{2\delta}{d(d-1)}+\tfrac{d-2}{d}$, for all $0\leq p\leq 1$ we have
$\mathbb{P}(\Psi\in\widehat{\mathcal{H}}\mid\Psi\not\in\mathcal{H})=o(1)\cdot s\,.$
We follow the approach in [BGP24]. Fix any $\Psi\subset[n]$ with $|\Psi|=d$, denote
$\mathcal{U}=\mathcal{U}(\mathcal{H})=\Big{\{}A\subset\Psi:|A|\geq 2,\exists\Phi\in\mathcal{H},A=\Phi\cap\Psi\Big{\}}\,.$
Then we have
$\mathbb{P}\big{(}\Psi\in\widehat{\mathcal{H}}\mid\Psi\not\in\mathcal{H}\big{)}=\sum_{\mathcal{U}:\Psi\not\in\mathcal{U}}\mathbb{P}\big{(}\Psi\in\widehat{\mathcal{H}};\mathcal{U}\big{)}=\sum_{\mathcal{U}:\Psi\not\in\mathcal{U}}\mathbb{P}(\mathcal{U})\mathbb{P}\big{(}\Psi\in\widehat{\mathcal{H}}\mid\mathcal{U}\big{)}\,.$
Denote $\mathcal{U}=\{A_{1},\ldots,A_{k}\}$. We then get from a simple union bound that
$\mathbb{P}(\mathcal{U})=\prod_{1\leq\ell\leq k}sn^{d-|A_{\ell}|}=\prod_{1\leq\ell\leq k}n^{1+\delta-|A_{\ell}|+o(1)}\,.$
Denoting $(x)_{+}\triangleq x\mathbf{1}_{x\geq 0}$ the positive part of $x$, we have
$\mathbb{P}\big{(}\Psi\in\widehat{\mathcal{H}}\mid\mathcal{U}\big{)}$ $=\mathbb{P}\big{(}(i,j)=1,\forall(i,j)\in E(\Psi)\setminus E(A_{\ell})\big{)}$
$\leq q^{(\binom{d}{2}-\sum_{1\leq\ell\leq k}\binom{|A_{\ell}|}{2})_{+}}=n^{(-1+\beta)(\binom{d}{2}-\sum_{1\leq\ell\leq k}\binom{|A_{\ell}|}{2})_{+}+o(1)}\,.$
Thus, we get that
$\mathbb{P}\big{(}\Psi\in\widehat{\mathcal{H}}\mid\Psi\not\in\mathcal{H}\big{)}\leq\sum_{\begin{subarray}{c}(A_{1},\ldots,A_{k})\\ A_{\ell}\subset\Psi,\Psi\neq A_{\ell}\end{subarray}}n^{-\omega(A_{1},\ldots,A_{k})+o(1)}\,,$
where
$\omega(A_{1},\ldots,A_{k})$ $=\sum_{1\leq\ell\leq k}\big{(}|A_{\ell}|-1-\delta\big{)}+(1-\beta)\Big{(}\tbinom{d}{2}-\sum_{1\leq i\leq k}\tbinom{|A_{\ell}|}{2}\Big{)}_{+}$
$>d-1-\delta+\sum_{1\leq\ell\leq k}\big{(}|A_{\ell}|-1-\delta\big{)}-\big{(}\tfrac{2}{d}-\tfrac{2\delta}{d(d-1)}\big{)}\sum_{1\leq\ell\leq k}\tbinom{|A_{i}|}{2}\,,$
where the inequality follows from $\beta<\tfrac{d-2}{d}+\tfrac{2\delta}{d(d-1)}$. Note that $12\leq|A_{\ell}|\leq d-$ 1 and $\delta<\tfrac{d-1}{d+1}$, we get that
$\sum_{1\leq\ell\leq k}\big{(}|A_{\ell}|-1-\delta\big{)}-\big{(}\tfrac{2}{d}-\tfrac{2\delta}{d(d-1)}\big{)}\sum_{1\leq\ell\leq k}\tbinom{|A_{\ell}|}{2}$
=\ = $-\tfrac{1}{d}\sum_{1\leq\ell\leq k}\big{(}|A_{\ell}|-d\big{)}\big{(}(1-\tfrac{\delta}{d-1})|A_{\ell}|-(1+\delta)\big{)}\geq 0\,,$
In conclusion, we obtain that $\omega(A_{1},\ldots,A_{k})>d-1-\delta$ for all $(A_{1},\ldots,A_{k})$. Since the possible choice of $(A_{1},\ldots,A_{k})$ is bounded by a constant depending only on $d$, we get that
$\mathbb{P}\big{(}\Psi\in\widehat{\mathcal{H}}\mid\Psi\not\in\mathcal{H}\big{)}\leq O(1)\cdot n^{-(d-1)+\delta-\Omega(1)}=o(1)\cdot s\,,$
as desired.
5 Proof of Proposition 1.8
We first prove Item (1). Define $\mu^{\prime}=\mu\big{(}\cdot\mid|\mathcal{H}|\leq(1+o(1))s\tbinom{n}{d}\big{)}$. Conditioned on $\mathcal{H}\sim\mu^{\prime}$, we define $\mathbb{P}^{\prime}$ by letting $\mathbb{P}^{\prime}(\cdot\mid\mathcal{H})\triangleq\mathbb{P}(\cdot\mid\mathcal{H})$. It is straightforward to check that $\operatorname{TV}(\mathbb{P}^{\prime},\mathbb{P})=o(1)$. Thus, by data processing inequality, it suffices to show that when $p=o(1),q=0$, for any estimator ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$ we have
(|^$\mathbb{P}^{\prime}\Big{(}|\widehat{\mathcal{H}}(A)\cap\mathcal{H}|\geq\delta\cdot s\tbinom{n}{d}\Big{)}=o(1)\mbox{ for any constant }\delta>0\,.$
Our proof relies on a generalization of Fano’s inequality which appears in e.g. [SW12]. The formulation we use here comes from [BHM12, Lemma 20] (see also [HM23, Theorem 3.1]). This generalization departs from the classical Fano’s inequality by adopting a broader notion of what constitutes a successful learning algorithm.
Lemma 5.1 (Lemma 20 in [BHM12]).
Let $\mathcal{W}$ be a hypothesis class of cardinality $M$ with $Z$ being a hypothesis drawn uniformly at random from $\mathcal{W}$. Let $X$ and $\widehat{Z}\in\mathcal{W}$ be random variables defined in such a way that $Z\longrightarrow X\longrightarrow\widehat{Z}$ constitutes a Markov chain. We define $d:\mathcal{W}\times\mathcal{W}\to\mathbb{R}_{\geq 0}$ to be a distance function and, for a given $d>0$, we say that a learner is successful if $d(Z,\widehat{Z})\leq d$. For any $h\in\mathcal{W}$, we further define $B_{d}(h)=\{h^{\prime}\in\mathcal{W}:d(h,h^{\prime})\leq d\}$ and we let and we let $M_{d}=\max_{h\in\mathcal{W}}\{|B_{d}(h)|\}$. Then we have
$\mathbb{P}\big{(}d(Z,\widehat{Z})\geq d\big{)}\geq 1-\tfrac{I(Z;X)+1}{\log(M/M_{d})}$
where $I(Z;X)$ is the mutual information between $Z$ and $X$.
For our purpose, for any $K=(1+o(1))s\tbinom{n}{d}$ we let $\mathcal{W}=\{\mathcal{H}\subset\tbinom{[n]}{d}:|\mathcal{H}|=K\}$. Noting that $\mathcal{H}\longrightarrow A\longrightarrow\widehat{\mathcal{H}}$ is a Markov chain and one can define a distance metric over $\mathcal{W}\times\mathcal{W}$ by taking $d(\mathcal{H},\mathcal{H}^{\prime})=\tfrac{|\mathcal{H}\triangle\mathcal{H}^{\prime}|}{2s\binom{n}{d}}$. For any constant $10\leq\epsilon<$ 1, we then have
${}\mathbb{P}\big{(}d(\mathcal{H},\widehat{\mathcal{H}})\leq\epsilon\mid|\mathcal{H}|=K\big{)}\leq\tfrac{I(\mathcal{H};A)+1}{\log(M/M_{\epsilon})}\,,$ (5.1)
where $M=|\mathcal{W}|$ and $M_{\epsilon}=|\{\mathcal{H}^{\prime}\in\mathcal{W}:|\mathcal{H}^{\prime}\cap\mathcal{H}|\geq(1-\epsilon)s\binom{n}{d}\}|$. Clearly we have (recall that $K=n^{1+\delta+o(1)}$)
$M=\binom{\tbinom{n}{d}}{K}=\exp\big{(}(d-1-\delta+o(1))K\log n\big{)}\,,$
$M_{\epsilon}\leq\binom{\tbinom{n}{d}}{\epsilon K}\cdot\binom{K}{(1-\epsilon)K}\leq\exp\big{(}(d-1-\delta+o(1))\epsilon K\log n\big{)}\,.$
We then have
${}\log(M/M_{\epsilon})\geq(1-\epsilon)(d-1-\delta-o(1))K\log n\,.$ (5.2)
In addition, since $p=o(1)$ and $q=0$, we then have
${}I(\mathcal{H};A)=H(A)-H(A\mid\mathcal{H})\leq H(A)\leq\log\Big{(}\binom{\binom{n}{2}}{pK\binom{d}{2}}\Big{)}\leq(1-\delta)pK\log n\,.$ (5.3)
Plugging (5.2) and (5.3) into (5.1), we get that
$\mathbb{P}\big{(}d(\mathcal{H},\widehat{\mathcal{H}})\leq\epsilon\mid|\mathcal{H}|=K\big{)}=o(1)\mbox{ for all }K=(1+o(1))s\tbinom{n}{d}\,.$
This verifies Item (1) of Proposition 1.8.
Now we turn to the proof of Item (2). Fix any $\iota=\Omega(1)$. The first step is to employ suitable truncations. Recall the definition of $\mathcal{E}$ in (2.3), (2.4) and (2.5). Note that
{${}\Big{\{}|\mathcal{H}|=(1+o(1))s\tbinom{n}{d}\Big{\}}\cap\Big{\{}|E(\mathcal{P}(\mathcal{J}))|\geq(1-o(1))\tbinom{d}{2}|\mathcal{J}|,\forall\mathcal{J}\subset\mathcal{H},|\mathcal{J}|\geq\iota s\tbinom{n}{d}\Big{\}}\subset\mathcal{E}\,,$ (5.4)
where the relation $\subset$ follows by taking $\mathcal{J}=\mathcal{H}$. Define $\mu^{\prime}=\mu\big{(}\mid\mathcal{E}\big{)}$. In addition, denote $\mathbb{P}_{\mathcal{H}}=\mathbb{P}(\cdot\mid\mathcal{H})$ and define
${}\mathbb{P}^{\prime}(A)=\mathbb{E}_{\mathcal{H}\sim\mu^{\prime}}\Big{[}\mathbb{P}_{\mathcal{H}}\Big{(}A\mid|E(A)|=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}\Big{)}\Big{]}\,.$ (5.5)
Lemma 5.2.
We have $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime})=o(1)$.
Proof.
Recall that in (2.6) we have shown that $\mu(\mathcal{E})=1-o(1)$ and thus $\operatorname{TV}(\mu,\mu^{\prime})=o(1)$. Fix any $\mathcal{H}$ such that $|\mathcal{H}|=(1+o(1))s\tbinom{n}{d}$ and $|E(\mathcal{P}(\mathcal{H}))|=(1-o(1))s\tbinom{n}{d}\tbinom{d}{2}$, it is easy to check that with probability $1-o(1)$ we have
$|E(A(\mathcal{H}))|=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}+O\Big{(}(q\tbinom{n}{2})^{\frac{1}{2}}\Big{)}=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}\,.$
Thus, we have $\operatorname{TV}(\mathbb{P}_{\mathcal{H}},\mathbb{P}^{\prime}_{\mathcal{H}})=o(1)$. Using
$\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime})\leq\operatorname{TV}(\mu,\mu^{\prime})+\mathbb{E}_{\mathcal{H}\sim\mu^{\prime}}\Big{[}\operatorname{TV}(\mathbb{P}_{\mathcal{H}},\mathbb{P}^{\prime}_{\mathcal{H}})\Big{]}\,,$
we see that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime})=o(1)$. ∎
To prove Item (2), without loss of generality, we assume $p=1$. A crucial observation (as shown in [BGPY25, Definition 1.6 and Lemma 2.5]) is that we only need to prove the following result: suppose we sample $(\mathcal{H},A)$ according to $\mathbb{P}^{\prime}$ and we sample $\mathcal{H}^{\prime}$ according to the posterior distribution of $\mathcal{H}$ given $A$. Then for all fixed $\iota=\Omega(1)$ we have
${}\mathbb{P}^{\prime}\Big{(}|\mathcal{H}\cap\mathcal{H}^{\prime}|\geq\iota s\tbinom{n}{d}\Big{)}=o(1)\,.$ (5.6)
We now focus on proving (5.6). By direct expansion we get that the left hand side of (5.6) equals
$\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime},A\\ |\mathcal{H}\cap\mathcal{H}^{\prime}|\geq\iota s\tbinom{n}{d}\end{subarray}}\frac{\mathbb{P}^{\prime}(\mathcal{H},A)\mathbb{P}^{\prime}(\mathcal{H}^{\prime},A)}{\mathbb{P}^{\prime}(A)}\,.$ (5.7)
Using Cauchy-Schwartz inequality, we see that (5.7) is bounded by
$\Bigg{(}\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime},A\\ |\mathcal{H}\cap\mathcal{H}^{\prime}|\geq\iota s\tbinom{n}{d}\end{subarray}}\frac{\mathbb{P}^{\prime}(\mathcal{H},A)\mathbb{P}^{\prime}(\mathcal{H}^{\prime},A)}{\mathbb{P}^{\prime}(A)^{2}}\Bigg{)}^{\frac{1}{2}}\Bigg{(}\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime},A\\ |\mathcal{H}\cap\mathcal{H}^{\prime}|\geq\iota s\tbinom{n}{d}\end{subarray}}\mathbb{P}^{\prime}(\mathcal{H},A)\mathbb{P}^{\prime}(\mathcal{H}^{\prime},A)\Bigg{)}^{\frac{1}{2}}\,.$ (5.8)
Note that (recall (5.5))
$\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime},A\\ |\mathcal{H}\cap\mathcal{H}^{\prime}|\geq\iota s\tbinom{n}{d}\end{subarray}}\frac{\mathbb{P}^{\prime}(\mathcal{H},A)\mathbb{P}^{\prime}(\mathcal{H}^{\prime},A)}{\mathbb{P}^{\prime}(A)^{2}}\leq\sum_{A:|E(A)|=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}}\sum_{\mathcal{H},\mathcal{H}^{\prime}}\mathbb{P}^{\prime}(\mathcal{H}\mid A)\mathbb{P}^{\prime}(\mathcal{H}^{\prime}\mid A)$
=\ = #$\#\Big{\{}A:|E(A)|=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}\Big{\}}$
[…]
where in the last inequality we use the fact that $q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}=(1+n^{-\Omega(1)})q\tbinom{n}{2}$, which follows from the assumption $\delta\leq\frac{d-1}{d+1}$. In addition, we have (recall (5.4))
$\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime},A\\ |\mathcal{H}\cap\mathcal{H}^{\prime}|\geq\iota s\tbinom{n}{d}\end{subarray}}\mathbb{P}^{\prime}(\mathcal{H},A)\mathbb{P}^{\prime}(\mathcal{H}^{\prime},A)=\sum_{I\geq\iota s\tbinom{n}{d}}\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime}\\ |\mathcal{H}\cap\mathcal{H}^{\prime}|=I\end{subarray}}\mu^{\prime}(\mathcal{H})\mu^{\prime}(\mathcal{H}^{\prime})\sum_{A}\mathbb{P}^{\prime}(A\mid\mathcal{H})\mathbb{P}^{\prime}(A\mid\mathcal{H}^{\prime})$
=\ = $\sum_{\begin{subarray}{c}I\geq\iota s\binom{n}{d}\\ J\geq 0\end{subarray}}\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime}:|\mathcal{H}\cap\mathcal{H}^{\prime}|=I\\ |E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))|=\binom{d}{2}I+J\end{subarray}}\mu^{\prime}(\mathcal{H})\mu^{\prime}(\mathcal{H}^{\prime})\sum_{\begin{subarray}{c}A:\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H}^{\prime})\subset A\\ |E(A)|=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}\end{subarray}}\mathbb{P}^{\prime}(A\mid\mathcal{H})\mathbb{P}^{\prime}(A\mid\mathcal{H}^{\prime})\,.$ (5.10)
Note that for all $\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H}^{\prime})\subset A$ and $|E(A)|=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}$, it holds that (recall that $q=n^{-1+\beta+o(1)}$)
$\mathbb{P}^{\prime}(A\mid\mathcal{H}),\mathbb{P}^{\prime}(A\mid\mathcal{H}^{\prime})$ $\leq q^{|E(A)|-(1+o(1))s\binom{d}{2}\binom{n}{d}}$
$=\exp\Big{(}-\big{(}q\tbinom{n}{2}+o(1)\cdot s\tbinom{n}{d}\tbinom{d}{2}\big{)}\log(q^{-1})\Big{)}\,.$
Thus, we get that (note that $|E(\mathcal{P}(\mathcal{H}))\cup E(\mathcal{P}(\mathcal{H}^{\prime}))|=(2+o(1))s\tbinom{n}{d}\tbinom{d}{2}-I\tbinom{d}{2}-J$)
$\sum_{\begin{subarray}{c}A:\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H}^{\prime})\subset A\\ |E(A)|=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}\end{subarray}}\mathbb{P}^{\prime}(A\mid\mathcal{H})\mathbb{P}^{\prime}(A\mid\mathcal{H}^{\prime})\leq\exp\Big{(}-2\big{(}q\tbinom{n}{2}+o(1)\cdot s\tbinom{n}{d}\tbinom{d}{2}\big{)}\log(q^{-1})\Big{)}\times$
#$\#\Big{\{}A:\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H}^{\prime})\subset A,|E(A)|=q\tbinom{n}{2}+(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}\Big{\}}$
[…]
In addition, we have
$\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime}:|\mathcal{H}\cap\mathcal{H}^{\prime}|=I\\ |E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))|=J+I\binom{d}{2}\end{subarray}}\mu^{\prime}(\mathcal{H})\mu^{\prime}(\mathcal{H}^{\prime})$
=\ = $[1+o(1)]\cdot\mathbb{P}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu\otimes\mu}\Big{(}|\mathcal{H}\cap\mathcal{H}^{\prime}|=I,|E(\mathcal{P}(\mathcal{H}))\cap E(\mathcal{P}(\mathcal{H}^{\prime}))|=I\tbinom{d}{2}+J\Big{)}\,.$ (5.12)
Recall that $s\tbinom{n}{d}=n^{1+\delta+o(1)}$, we have
$\mathbb{P}_{\mathcal{H},\mathcal{H}^{\prime}\sim\mu\otimes\mu}\Big{(}|\mathcal{H}\cap\mathcal{H}^{\prime}|=I\Big{)}\leq$ $\binom{\binom{n}{d}}{(1+o(1))s\binom{n}{d}-I}\binom{(1+o(1))s\binom{n}{d}}{I}\Bigg{/}\binom{\binom{n}{d}}{(1+o(1))s\binom{n}{d}}$
[…]
In addition, given $|\mathcal{J}|=I$ with $\mathcal{H}\cap\mathcal{H}^{\prime}=\mathcal{J}$, we have (denote $\mathcal{K}=\mathcal{H}\setminus\mathcal{J}$ and $\mathcal{K}^{\prime}=\mathcal{H}^{\prime}\setminus\mathcal{J}$)
$(\mathcal{K},\mathcal{K}^{\prime})\sim\widetilde{\mu}\otimes\widetilde{\mu}\Big{(}\cdot\mid\mathcal{K}\cap\mathcal{K}^{\prime}=\emptyset\Big{)}\,,$
where $\widetilde{\mu}$ is the distribution such that each element in $\operatorname{U}\setminus\mathcal{J}$ is chosen independently with probability $s$. Note that
$\mathbb{P}_{(\mathcal{K},\mathcal{K}^{\prime})\sim\widetilde{\mu}\otimes\widetilde{\mu}}\Big{(}\mathcal{K}\cap\mathcal{K}^{\prime}=\emptyset\Big{)}=(1-s^{2})^{\binom{n}{d}}\geq\exp\big{(}-s^{2}\tbinom{d}{2}\big{)}=\exp\big{(}o(1)\cdot I\big{)}\,,$
where the last equality follows from $I\geq\iota s\binom{n}{d}$. In addition, similarly as in Lemma 3.3, we get that
$\mathbb{P}_{(\mathcal{K},\mathcal{K}^{\prime})\sim\widetilde{\mu}\otimes\widetilde{\mu}}\Big{(}|E(\mathcal{K})\cap E(\mathcal{K}^{\prime})|\geq J\Big{)}\leq\exp\Big{(}-J^{2}/n^{2\delta+o(1)}\Big{)}\,.$ (5.14)
Thus, combining (5.13) and (5.14) we have
(5.12)$\leq\exp\Big{(}-(1+o(1))I\log(s^{-1})-J^{2}/n^{2\delta+o(1)}\Big{)}\,.$ ) . (5.15)
Combining (5.11), (5.12) and (5.15), we get that (5.10) is bounded by
$\sum_{\begin{subarray}{c}I\geq\iota s\binom{n}{d}\\ J\geq 0\end{subarray}}\exp\Big{(}-q\tbinom{n}{2}\log(q^{-1})-(1+o(1))(s\tbinom{n}{d}\tbinom{d}{2}-I\tbinom{d}{2}-J)\log(q^{-1})\Big{)}$
$\cdot\exp\Big{(}-(1+o(1))I\log(s^{-1})-J^{2}/n^{2\delta+o(1)}\Big{)}$
[…]
where in the first inequality we use that $I\geq\Omega(1)\cdot s\tbinom{n}{d}=n^{1+\delta+o(1)}$ and
$\sum_{J\geq 0}\exp\Big{(}J\log(q^{-1})-J^{2}/n^{2\delta+o(1)}\Big{)}=\exp\Big{(}n^{2\delta+o(1)}\Big{)}=\exp\Big{(}o(1)\cdot I\Big{)}\,,$
and in the last inequality we use the fact that
$\tbinom{d}{2}\log(q^{-1})-\log(s^{-1})$ $=\tbinom{d}{2}(1-\beta+o(1))-(d-1-\delta+o(1))$
$=\Omega(1)\quad\mbox{ when }\beta>\tfrac{d-2}{d}+\tfrac{2\delta}{d(d-1)}\,.$
Plugging (5.16) and (5.9) into (5.8), we obtain that (5.7) is bounded by
$\exp\Big{(}-\big{(}q\tbinom{n}{2}-(1+\Omega(1))s\tbinom{n}{d}\tbinom{d}{2}\big{)}\log(q^{-1})\Big{)}\cdot\exp\Big{(}\big{(}q\tbinom{n}{2}-(1+o(1))s\tbinom{n}{d}\tbinom{d}{2}\big{)}\log(q^{-1})\Big{)}=o(1)\,.$
This proves (5.6).
6 Proof of Proposition 1.9
In this section, without loss of generality, we assume $q=0$. Following the approach in [DWXY23], we only need to prove the following result: suppose we sample $(\mathcal{H},A)$ according to $\mathbb{P}$ and we sample $\mathcal{H}^{\prime}$ according to the posterior distribution of $\mathcal{H}$ given $A$. Then we have
${}\mathbb{P}\Big{(}|\mathcal{H}\triangle\mathcal{H}^{\prime}|\leq o(1)\cdot s\tbinom{n}{d}\Big{)}=o(1)\,.$ (6.1)
To see this, note that for any estimator ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(A)$, we have that $(\widehat{\mathcal{H}},\mathcal{H})$ and $(\widehat{\mathcal{H}},\mathcal{H}^{\prime})$ are equal in law. Thus, we have
$o(1)=\mathbb{P}\Big{(}|\mathcal{H}\triangle\mathcal{H}^{\prime}|\leq o(1)\cdot s\tbinom{n}{d}\Big{)}$ (|^$\geq\mathbb{P}\Big{(}|\mathcal{H}^{\prime}\triangle\widehat{\mathcal{H}}|\leq o(1)\cdot s\tbinom{n}{d}\Big{)}+\mathbb{P}\Big{(}|\mathcal{H}\triangle\widehat{\mathcal{H}}|\leq o(1)\cdot s\tbinom{n}{d}\Big{)}$
$=2\mathbb{P}\Big{(}|\mathcal{H}\triangle\widehat{\mathcal{H}}|\leq o(1)\cdot s\tbinom{n}{d}\Big{)}\,,$
where the inequality follows from the fact that
$|\mathcal{H}\triangle\widehat{\mathcal{H}}|+|\mathcal{H}^{\prime}\triangle\widehat{\mathcal{H}}|\geq|\mathcal{H}\triangle\mathcal{H}^{\prime}|\,.$
This yields Proposition 1.9. We now prove (6.1). For any $\mathcal{H}\in\{0,1\}^{\binom{[n]}{d}}$ and $A\in\{0,1\}^{\binom{[n]}{2}}$, define
${}\mathsf{Empty}(\mathcal{H};A)\triangleq\Big{\{}\Psi\in\mathcal{H}:E(\mathcal{P}(\Psi))\cap E(A)=\emptyset\Big{\}}\,.$ (6.2)
Define
[…]
$|E(\mathcal{P}(\mathsf{Empty}(\mathcal{H};A)))|\geq(1-o(1))\tbinom{d}{2}|\mathsf{Empty}(\mathcal{H};A)|\Big{\}}\,.$
It is easy to check that $\mathbb{P}((\mathcal{H},A)\in\mathcal{G})=1-o(1)$. In addition, for any $(\mathcal{H},A)\in\mathcal{G}$, denote
[…]
It is straightforward to check that for $(\mathcal{H},A)\in\mathcal{G}$
#$\#\mathsf{Blur}(\mathcal{H};A)\geq\binom{\binom{n}{d}-(1+o(1))s\binom{n}{d}}{|\mathsf{Empty}(\mathcal{H};A)|}\geq\exp\Big{(}\Omega(1)\cdot s\tbinom{n}{d}\log n\Big{)}\,.$
In addition, for all $\mathcal{H}^{\prime}\in\mathsf{Blur}(\mathcal{H};A)$ we have
$\frac{\mathbb{P}(\mathcal{H}\mid A)}{\mathbb{P}(\mathcal{H}^{\prime}\mid A)}$ $=\frac{\mathbb{P}(\mathcal{H},A)}{\mathbb{P}(\mathcal{H}^{\prime},A)}=\frac{\mu(\mathcal{H})\mathbb{P}_{\mathcal{H}}(A)}{\mu(\mathcal{H}^{\prime})\mathbb{P}_{\mathcal{H}^{\prime}}(A)}=\frac{\mathbb{P}_{\mathcal{H}}(A)}{\mathbb{P}_{\mathcal{H}^{\prime}}(A)}$
$=\frac{p^{|E(A)|}(1-p)^{|E(\mathcal{P}(\mathcal{H}))\setminus E(A)|}}{p^{|E(A)|}(1-p)^{|E(\mathcal{P}(\mathcal{H}^{\prime}))\setminus E(A)|}}\leq\exp\Big{(}o(1)\cdot s\tbinom{n}{d}\Big{)}\,,$
where the third equality follows from $|\mathcal{H}|=|\mathcal{H}^{\prime}|$ and the inequality follows from
$|E(\mathcal{P}(\mathcal{H}))\setminus E(A)|-|E(\mathcal{P}(\mathcal{H}^{\prime}))\setminus E(A)|$
=\ = $|E(\mathcal{P}(\mathsf{Empty}(\mathcal{H};A)))|-|E(\mathcal{P}(\mathcal{H}^{\prime}\cap\mathsf{Empty}(\mathcal{H};A)))\setminus E(A)|$
[…]
Thus, we see that for $(\mathcal{H},A)\in\mathcal{G}$
${}\mathbb{P}(\mathcal{H}\mid A)\leq\frac{\exp(o(1)\cdot s\tbinom{n}{d})}{|\mathsf{Blur}(\mathcal{H};A)|}=\exp\Big{(}-\Omega(1)\cdot s\tbinom{n}{d}\log n\Big{)}\,.$ (6.5)
Now we return to the proof of (6.1). Clearly we can write the left hand side of (6.1) as
$\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime},A\\ |\mathcal{H}\triangle\mathcal{H}^{\prime}|\leq o(1)\cdot s\binom{n}{d}\end{subarray}}\frac{\mathbb{P}(\mathcal{H},A)\mathbb{P}(\mathcal{H}^{\prime},A)}{\mathbb{P}(A)}$
[…]
Note that
$\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime};A\\ (\mathcal{H},A)\not\in\mathcal{G}\end{subarray}}\frac{\mathbb{P}(\mathcal{H},A)\mathbb{P}(\mathcal{H}^{\prime},A)}{\mathbb{P}(A)}=\sum_{\begin{subarray}{c}(\mathcal{H},A)\not\in\mathcal{G}\end{subarray}}\mathbb{P}(\mathcal{H},A)\sum_{\mathcal{H}^{\prime}}\mathbb{P}(\mathcal{H}^{\prime}\mid A)\leq\mathbb{P}((\mathcal{H},A)\not\in\mathcal{G})=o(1)\,.$ (6.7)
In addition, we have
$\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime};A:(\mathcal{H},A),(\mathcal{H}^{\prime},A)\in\mathcal{G}\\ |\mathcal{H}\triangle\mathcal{H}^{\prime}|\leq o(1)\cdot s\binom{n}{d}\end{subarray}}\frac{\mathbb{P}(\mathcal{H},A)\mathbb{P}(\mathcal{H}^{\prime},A)}{\mathbb{P}(A)}=\sum_{\begin{subarray}{c}\mathcal{H},\mathcal{H}^{\prime};A:(\mathcal{H},A),(\mathcal{H}^{\prime},A)\in\mathcal{G}\\ |\mathcal{H}\triangle\mathcal{H}^{\prime}|\leq o(1)\cdot s\binom{n}{d}\end{subarray}}\mathbb{P}(\mathcal{H}^{\prime},A)\mathbb{P}(\mathcal{H}\mid A)$
[…]
where the third inequality follows from for $|\mathcal{H}^{\prime}|=(1+o(1))s\tbinom{n}{d}$
#$\#\Big{\{}\mathcal{H}:|\mathcal{H}\triangle\mathcal{H}^{\prime}|\leq o(1)\cdot s\tbinom{n}{d}\Big{\}}\leq 2^{|\mathcal{H}^{\prime}|}\cdot\binom{\binom{n}{d}}{o(1)\cdot s\binom{n}{d}}=\exp\Big{(}o(1)\cdot s\tbinom{n}{d}\log n\Big{)}\,.$
And the result follows directly by plugging (6.7) and (6.8) into (6.6).
Acknowledgment
S. Gong and Z. Li are partially supported by National Key R&\&&D Program of China (Project No. 2023YFA1010103) and NSFC Key Program (Project No.12231002).
Appendix A Detection with a modified null distribution
In the main text, we have cited the presence of a detection-recovery gap as a source of difficulty for obtaining tight thresholds for recovery. One might wonder whether this can be remedied simply by choosing a better null distribution $\mathbb{Q}$ that is harder to distinguish from the planted distribution $\mathbb{P}$. In this section, we will argue that this approach does not work. Recall Definition 1.1. Define ~=~$\widetilde{\mathbb{Q}}=\widetilde{\mathbb{Q}}_{n}$ to be the Erdős-Rényi graph with edge-density
${}\widetilde{q}=\tfrac{1}{\binom{n}{2}}\Big{(}(p-q)s\tbinom{n}{d}\tbinom{d}{2}\Big{)}+q\,.$ (A.1)
Clearly $\widetilde{\mathbb{Q}}$ and $\mathbb{P}$ have the same expectational edge-density. However, we will show that it is still possible to distinguish $\mathbb{P}$ from $\widetilde{\mathbb{Q}}$ when $\delta,\beta<\tfrac{d-1}{d+1}$. Consider the “clique counting” statistic
${}g(A)=\sum_{\Psi\in\operatorname{U}}\mathbf{1}_{\{\mathcal{P}(\Psi)\subset A\}}\,.$ (A.2)
Proposition A.1.
Suppose $\delta,\beta<\tfrac{d-1}{d+1}$. We then have
$\mathbb{P}\Big{(}g(A)\geq q^{\binom{d}{2}}\tbinom{n}{d}+\tfrac{1}{2}ps\tbinom{n}{d}\Big{)}=1-o(1)\,,$
~$\widetilde{\mathbb{Q}}\Big{(}g(A)<q^{\binom{d}{2}}\tbinom{n}{d}+\tfrac{1}{2}ps\tbinom{n}{d}\Big{)}=1-o(1)\,.$
Proof.
Denote $\kappa=\max\{\delta,\beta\}$ and recall from (A.1) that $\widetilde{q}=n^{-1+\kappa+o(1)}$. When $\kappa<\tfrac{d-3}{d-1}$, it can be easily checked using Markov’s inequality that ~$\widetilde{\mathbb{Q}}(g(A)=0)=1-o(1)$. When $\kappa\geq\tfrac{d-3}{d-1}$, we have
$\operatorname{Var}_{\widetilde{\mathbb{Q}}}(g(A))=\sum_{\Psi,\Phi\in\operatorname{U}}\operatorname{Cov}\Big{(}\mathbf{1}_{\{\mathcal{P}(\Psi)\subset A\}},\mathbf{1}_{\{\mathcal{P}(\Phi)\subset A\}}\Big{)}\,.$
Note that for $2\leq m\leq d$, we have #$\#\{\Psi,\Phi:|V(\Psi)\cap V(\Phi)|=m\}=n^{2d-m+o(1)}$ and for $|V(\Psi)\cap V(\Phi)|=m$ we have
$\operatorname{Cov}\big{(}\mathbf{1}_{\{\mathcal{P}(\Psi)\subset A\}},\mathbf{1}_{\{\mathcal{P}(\Phi)\subset A\}}\big{)}=\widetilde{q}^{2\binom{d}{2}-\binom{m}{2}}\,.$
Thus, we have
$\operatorname{Var}_{\widetilde{\mathbb{Q}}}(g(A))=\mathbb{E}_{\widetilde{\mathbb{Q}}}[g(A)]^{2}\cdot\sum_{2\leq m\leq d}n^{-m+(1-\kappa)\binom{m}{2}+o(1)}\,.$
Note that when $\kappa>\tfrac{d-3}{d-1}$, we have $-m+(1-\kappa)\binom{m}{2}$ is monotone decreasing in $m$ when $2\leq m\leq d$. Thus,
$\operatorname{Var}_{\widetilde{\mathbb{Q}}}(g(A))=\mathbb{E}_{\widetilde{\mathbb{Q}}}[g(A)]^{2}\cdot n^{-1-\kappa+o(1)}\,.$
This yields that
$\operatorname{Var}_{\widetilde{\mathbb{Q}}}(g(A))^{\tfrac{1}{2}}\leq n^{d-\binom{d}{2}(1-\kappa)-\tfrac{1+\kappa}{2}+o(1)}\leq n^{1+\delta-\Omega(1)}\,,$
where the last inequality follows from our assumption that
$d-\tbinom{d}{2}(1-\kappa)-\tfrac{1+\kappa}{2}=1+\delta-\Omega(1)\mbox{ when }\beta,\delta<\tfrac{d-1}{d+1}\,.$
In addition, we have
$\mathbb{E}_{\widetilde{\mathbb{Q}}}[g(A)]=\widetilde{q}^{\binom{d}{2}}\tbinom{n}{d}=q^{\binom{d}{2}}\tbinom{n}{d}+O(1)\cdot\tbinom{n}{d}\sum_{k=0}^{\binom{d}{2}-1}q^{k}(\widetilde{q}-q)^{\binom{d}{2}-k}$
Note that $\widetilde{q}-q=n^{-1+\delta+o(1)}$ and $q=n^{-1+\beta+o(1)}$, thus we have for all $10\leq k\leq\tbinom{d}{2}-$ 1
$\tbinom{n}{d}q^{k}(\widetilde{q}-q)^{\binom{d}{2}-k}$ $=n^{d-k(1-\beta)-(\binom{d}{2}-k)(1-\delta)+o(1)}$
$\leq n^{d-(\tbinom{d}{2}-1)(1-\kappa)-(1-\delta)+o(1)}\ll ps\tbinom{n}{d}=n^{(1+\delta)+o(1)}\,,$
where the last inequality follows from the fact that
$d-(\tbinom{d}{2}-1)(1-\kappa)-(1-\delta)<(1+\delta)\mbox{ when }\delta,\beta<\tfrac{d-1}{d+1}\,.$
In conclusion, we have shown that
$\mathbb{E}_{\widetilde{\mathbb{Q}}}[g(A)]=q^{\binom{d}{2}}\tbinom{n}{d}+o(1)\cdot ps\tbinom{n}{d}\mbox{ and }\operatorname{Var}_{\widetilde{\mathbb{Q}}}[g(A)]^{\frac{1}{2}}=o(1)\cdot ps\tbinom{n}{d}\,.$
Thus from standard Chebyshev’s inequality we have
~$\widetilde{\mathbb{Q}}\Big{(}g(A)<q^{\binom{d}{2}}\tbinom{n}{d}+\tfrac{1}{2}ps\tbinom{n}{d}\Big{)}=1-o(1)\,.$
In addition, note that under $\mathbb{P}$ we have $|\mathcal{H}|\geq 0.9ps\tbinom{n}{d}$ with probability $1-o(1)$. Thus, it is clear that
$\mathbb{P}\Big{(}g(A)\geq q^{\binom{d}{2}}\tbinom{n}{d}+\tfrac{1}{2}ps\tbinom{n}{d}\Big{)}=1-o(1)\,.\qed$